OpenStudy (anonymous):

Which of the following represents inverse variation? (A) y/x = -1/2 (B) y = -x (C) y = 1 - x (D) y = 7/x

OpenStudy (mertsj):

If an equation can be written xy = a constant, then it is inverse variation.

OpenStudy (mertsj):

Which one can be written that way?

OpenStudy (anonymous):

what mertsj said. the bigger x, the smaller y

OpenStudy (anonymous):

So wich one is it?

OpenStudy (anonymous):

Wait is it a?

OpenStudy (amistre64):

xy=a when we solve for y we get?

OpenStudy (amistre64):

xy = a /x /x ------- y = a/x